Solve using the elimination method.
4x-8y=48
11x+3y=-105.5

Answer by ewatrrr(24785)

 
Hi,
Solving using the elimination method.
4x-8y=48
11x+3y=-105.5
12/8x-3y = 18 |Multiplying 1st EQ by 3/8 to eliminate y
11x +3y= -105.5
100/8 x = -87.5
x = -87.5 * 8/100
x = -7 and y = -9.5 [(-28-48)/8 = -9.5]
CHECKING our Answer***
11*-7 + 3*-9.5 = -77 + 28.5 = -105.5